Grade 11General Physics

A man swimming downstream overcome a float at a point M.After travelling distance D he turned back and passed the float at a distance of D/2 from the point M,then the ratio of speed of swimmer with respect to still water to the speed of the river will be?

surya
10 Years agoGrade 11

1 Answer

parth pandey
6 Years ago
velocity of man =Vm
velocity of water=Vw
time taken for travelling distance D after point M by swimmer t = D/Vm+Vw    (velocity is added as its moving downstream)
 time taken for travelling D/2 after taking turn t= D/2/Vm-Vw ,  t=D/2(Vm-Vw)
TIME TAKEN BY FLOAT TO TRAVEL D/2 =D/2Vw
now time taken by swimmer to travel D andD/2 = time taken by float to travel D/2
SO,
d/Vm+Vw +D/2(Vm-Vw)= D/2Vw
now after solving we get
Vm/Vw = 3